OpenStudy (anonymous):

Factor this expression.. 3h2+32h+20..... help

OpenStudy (anonymous):

you mean 3h^2+32h+20?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

multiply 20 by the coefficient of h^2 which is 3 first. so you'll have h^2 + 32h +60, then you can factor that to (h+30)(h+2) then divide each of those answers by the original coefficient of 3. so h+30/3 = (h+10) and (h+2)/3 = (3h+2, because dividing 2 by 3 gives you a fraction, you bring the denominator of that fraction to the variable. so (h+10) and (3h+2) are your answers..

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

if you ever have a coefficient, and you can't take a common coefficient out, you do that. just multiply the coefficient of x^2 to the number without a variable at the end, factor it normally, then divide the 2 factors by the number you multiplied the last number by in the beginning.

OpenStudy (anonymous):

oh kk

OpenStudy (anonymous):

hope that helps even a little bit.. :)

OpenStudy (anonymous):

thanks